OpenStudy (vampirediaries):

A model rocket flies horizontally off the edge of a cliff at a velocity of 50.0 m/s. If the canyon below is 100.0 m deep, how far in the horizontal direction does the rocket travel from the edge of the cliff? (Assume no air resistance and that a^y=g=9.81 m/s^2.) A. 226 m B. 245 m C. 350 m D. 526 m

OpenStudy (anonymous):

Step 1: Find the time taken for the projectile to hit the ground. The only force acting the on the rocket is gravity. Therefore, we calculate the time through the free-fall equation. y = 0.5 * g * t^2 100 = 0.5 * 9.81 * t^2 t^2 = 20.39 t = 4.52 s So the rocket will be in the air for 4.52 s. Step 2: Find how far the rocket travels horizontally, knowing it is off the ground for 4.52 s. x = v * t = 50.0 * 4.52 = 226 m The rocket will travel 226 m in the horizontal direction. The answer is A.